OpenStudy (anonymous):

Choose the correct simplification of the expression the quantity a to the 4th power over b to the 2nd power all raised to the 3rd power. a to the 7th power over b to the 5th power a to the 12th power over b to the 6th power ab9 a12b6

OpenStudy (anonymous):

Here is what I think it is \[ \left(\frac{a^4}{b^2}\right)^3 = \frac{a^{4*3}}{b^{2*3}}=\frac{a^{12}}{b^6}\]

OpenStudy (anonymous):

ahh thank youu

OpenStudy (anonymous):

can you help with another?

OpenStudy (anonymous):

ya sure

OpenStudy (anonymous):

i shldnt be giving you away answers but ....

OpenStudy (anonymous):

hahah its one of my last assignments before i finish so I'm trying to just get it over with

OpenStudy (anonymous):

Ok go for it

OpenStudy (anonymous):

Graph 1 shows the sale of jackets at Harry's store for different values of average monthly temperature. Graph 2 shows some additional data points: graph 1 Plot ordered pairs 0, 100 and 2, 98 and 4, 96 and 5, 95 and 8, 92 and 10, 88 and 15, 84 and 18, 80 and 20, 80 and 23, 77 and 25, 75. Draw the line of <strong>best</strong> fit, graph 2 Plot ordered pairs 6, 90 and 7, 82 and 12, 79 and 13, 78 and 15, 76 and 18, 70 and 22, 69 and 28, 65 and 27, 65 and 28, 62 If the points shown on graph 2 are included on graph 1, how will the line of best fit for graph 1 be affected? It will shift up slightly. It will remain the same. It will shift down slightly. It will shift up considerably.

OpenStudy (anonymous):

It will shift down slightly... Since as you can see graph B is slightly lower so when you mix the two graphs together the line of best fit will shift slightly lower

OpenStudy (anonymous):

i appreciate this soooo much